# HLAT: High-quality Large Language Model Pre-trained on AWS Trainium

### v1

@@ Abstract.
Getting large language models (LLMs) to perform well on the downstream tasks requires pre-training over trillions of tokens. This typically demands a large number of powerful computational devices in addition to a stable distributed training framework to accelerate the training. The growing number of applications leveraging AI/ML had led to a scarcity of the expensive conventional accelerators (such as GPUs), which begs the need for the alternative specialized-accelerators that are scalable and cost-efficient. AWS Trainium  is the second-generation machine learning accelerator that has been purposely built for training large deep learning models. Its corresponding instance, Amazon EC2 trn1, is an alternative to GPU instances for LLM training. However, training LLMs with billions of parameters on trn1 is challenging due to its relatively nascent software ecosystem. In this paper, we showcase HLAT: a 7 billion parameter decoder-only LLM pre-trained using trn1 instances over 1.8 trillion tokens. The performance of HLAT is benchmarked against popular open source baseline models including LLaMA and OpenLLaMA, which have been trained on NVIDIA GPUs and Google TPUs, respectively. On various evaluation tasks, we show that HLAT achieves model quality on par with the baselines. We also share the best practice of using the Neuron Distributed Training Library (NDTL), a customized distributed training library for AWS Trainium  to achieve efficient training. Our work demonstrates that AWS Trainium  powered by the NDTL is able to successfully pre-train state-of-the-art LLM models with high performance and cost-effectiveness.
LLMs, pre-training, AWS Trainium, distributed training
Both authors contributed equally to this research.
conference: ; ;
1. Introduction
@@
Large language models (LLMs), based on transformer architecture (Vaswani et al., 2017) and trained on massive text data, is the most recent breakthrough in artificial intelligence. They not only show remarkable capabilities in understanding and generating text (Li et al., 2022), but offer immense potential across diverse downstream tasks, such as machine translation (Hendy et al., 2023), information retrieval (Zhu et al., 2023), code generation (Roziere et al., 2023) and so on (Zhao et al., 2023b).
@@
Pre-training is the crucial first step in building LLMs because it lays the foundation for their impressive capabilities. It initializes the model with random weights, and trains the model to convergence using tokens from a large text corpus. The training process is designed to be self-supervised. For decoder-only model, such as GPT (Brown et al., 2020) and LLaMA (Touvron et al., 2023a), the model is trained to predict the next token given a sequence of previous tokens. Eventually, the model learns everything ranging from syntax and semantics to world knowledge and commonsense reasoning with a large amount of training data. Pre-training provides the raw material - the language skills and understanding, which facilitate the subsequent fine tuning from different downstream tasks.
@@
Since pre-training requires a large amount of training data (trillions of tokens), it demands highly on computational resources. Advanced AI accelerators, such as AWS Trainium  [^1] https://aws.amazon.com/machine-learning/trainium, Google TPU  [^2] https://cloud.google.com/tpu, and NVIDIA A100/H100 GPUs  [^3] https://www.nvidia.com/en-us/data-center/a100;https://www.nvidia.com/en-us/data-center/h100, have been specifically designed for this kind of workloads. These AI accelerators are often integrated with dedicated tensor processing units which offer fast matrix operations and high training throughput. They also have much larger on-chip memory (tens of GBs per accelerator) and high communication bandwidth (hundreds of Gbps) between accelerators across different machines, which allows pre-training of larger models with efficient hardware utilization.
@@
Even with the powerful AI accelerators, due to the sheer size and complexity of LLMs, it’s impractical to train them on a single device. In other words, a single accelerator simply doesn’t have enough memory or processing power to handle the massive datasets, model parameters, and intricate calculations involved in LLM training. Practitioners rely on distributed training libraries (Zhao et al., 2023b) to orchestrate a number of accelerators to conduct the training together. Distributed libraries can shard the model parameters and optimizer states across multiple accelerators with different kinds of parallelism strategy. They also spread the workload across multiple machines, effectively tapping into a combined pool of resources, allowing to train the models at the scale of multi-billions of parameters. Additionally, by splitting the workload, distributed libraries significantly reduce the training time.
@@
Although there have been many successful demonstrations of pre-training LLMs on conventional accelerators (GPUs and TPUs) using state-of-the-art distributed training libraries (Rasley et al., 2020; Shoeybi et al., 2020; FairScale authors, 2021; Zhao et al., 2023a), training LLMs with billions of parameters on AWS Trainium  is still challenging. First, Trainium  uses a relatively nascent software ecosystem ranging from runtime, compiler, to distributed training library. The training script developed for other accelerators needs to be adjusted to comply with the low-level APIs and operators supported by Trainium. Second, the optimal training configurations that ensure stable convergence and optimal training throughput may also differ from other accelerators, e.g., level of precision, dimensions of 3D parallelism, compiler flags to mention a few. On the other hand, Amazon EC2 trn1 instance, equipped with AWS Trainium  accelerators, provides the comparable computation power to Amazon EC2 p4d instance, equipped with Nvidia A100 40GB GPUs, but comes with only  60% of the price (details in Section 2). This makes it appealing to fully utilize the compute power of AWS Trainium  for LLM pre-training.
In this paper, we for the first time provide an end-to-end LLM pre-training practice on top of AWS Trainium, demonstrating the effectiveness and efficiency of this accelerator. Specifically, we make the following contributions:
@@
  • We pre-train HLAT (High-quality LLM pre-trained on AWS Trainium), a 7B model following the architecture described in (Touvron et al., 2023a, b) from scratch using a total training budget of 1.8 trillion tokens. The pre-training is performed up to 64 Amazon EC2 trn1.32xlarge instances with totalling up to 1024 AWS Trainium  accelerators.
@@
  • We evaluate the pre-trained models on various evaluation tasks including commonsense reasoning, world knowledge, math, coding, etc. The results show that our pre-trained model provides comparable performance with 7B models trained on other AI accelerators and distributed training frameworks, including LLaMA-1, LLaMA-2, OpenLLaMA-1, and OpenLLaMA-2. We also evaluate the intermediate checkpoints during the training process for additional insights.
  • We provide some best practices of pre-training process, e.g., different sharding strategies, training precisions, and fault tolerance mechanism, on AWS Trainium  and Neuron Distributed Training Library (NDTL)  [^4] https://awsdocs-neuron.readthedocs-hosted.com/en/latest/libraries/neuronx-distributed/index.html. In addition, we design a novel dataloader for online sample packing which performs both tokenization and example packing during training. For pretraining on large datasets, our dataloader saves significant developer time and compute resources.
2. Background - Distributed Training on Trainium
@@
AWS Trainium is the second-generation machine learning accelerator that AWS purposely built for deep learning training. Each Trainium  accelerator includes two NeuronCores. Each NeuronCore has 16 GB of high-bandwidth memory, and delivers up to 95 TFLOPS of FP16/BF16 compute power. In this study, we trained our model on Amazon EC2 trn1.32xlarge instances: each instance is equipped with 16 Trainium  accelerators, and supports 800 Gbps intra-instance network bandwith through NeuronLink. The aggregating compute power of Amazon EC2 trn1.32xlarge is 3040 TFLOPS in FP16/BF16, slightly higher to its GPU instance counterpart Amazon EC2 p4d.24xlarge at 2496 TFLOPS, but at a much lower price (on demand hourly rate: trn1.32xlarge \$21.50 vs. p4d.24xlarge \$32.77).
@@
AWS Neuron is a software development kit (SDK) [^5] https://github.com/aws-neuron/aws-neuron-sdk with a compiler, runtime, and profiling tools that unlocks high-performance and cost-effective deep learning acceleration on AWS Trainium. Neuron is natively integrated with PyTorch (Zhao et al., 2023a) and TensorFlow (Abadi et al., 2015), and offers features such as FP32 autocasting, stochastic rounding, collective communication, custom operators, and so on.
Neuron Distributed Training Library (NDTL), as part of Neuron SDK, is developed to support high-efficiency distributed training on Trainium:
  • NDTL supports a variety of existing distributed training techniques, such as 3D parallelism (Shoeybi et al., 2020) , i.e., Tensor Parallelism (TP), Pipeline Parallelism (PP) and Data Parallelism (DP). To reduce the activation memory during training, activation checkpointing (Chen et al., 2016) and sequence parallelism (Korthikanti et al., 2022) are naturally supported with the 3D parallelism. NDTL also supports Zero Redundancy Optimizer Stage 1 (ZeRO-1) (Rajbhandari et al., 2020) to shard optimizer states. 3D parallelism and ZeRO-1 can be applied simultaneously during training.
  • NDTL provides unified interfaces to port users’ models and training scripts on Trainium  accelerators. The NDTL interfaces are friendly to Huggingface transformers library (Wolf et al., 2020). To train models from Huggingface transformers on Trainium  accelerators with NDTL, it only requires simple code changes in certain layers of the model.
  • NDTL supports TP with mixed degrees, i.e., users can use more than one TP degrees to shard different model parameters. TP with mixed degrees is used to handle the case when the LLM has some model parts are not compatible with a unified large TP degree, e.g., when the Grouped Query Attention (GQA) (Hudson and Manning, 2019) is applied, which prevents for using a unified TP degree to the whole model.
  • NDTL supports automatic fault recovery and checkpointing. Checkpointing is optimized to save/load on different machines quickly at the same time, and even able to save asynchronously. In case of hardware failures or communication timeouts, NDTL can automatically restart training from latest saved checkpoints without manual intervention, which is critical for maintaining system uptime and stability.
3. Method
3.1. Model Architecture
@@
HLAT  adopts the decoder-only transformer architecture and applies same modifications used in LLaMA 7B (Touvron et al., 2023a, b), including pre-normalization with RMSNorm, SwiGLU activation function, and Rotary Embeddings. The model is trained with a maximum sequence length of 4096.
3.2. Training Hyperparameters
@@
We adopt same training hyperparameters as LLaMA 7B (Touvron et al., 2023a, b) models. Specifically, we use a cosine learning rate scheduler with maximum learning rate of $3e^{-4}$ and minimum learning rate of $3e^{-5}$. We use a linear warmup of 2000 steps. The overall learning rate scheduler is plotted in Figure 1. We use AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. We use weight decay value of 0.1 for all parameters, including normalization weights. Gradient-norm clipping of 1.0 is applied for training stability.
We use 64 nodes with tensor parallel degree of 8, pipeline parallel degree of 1, and data parallel degree of 256. We use global batch size of 1024 sequences with maximum sequence length of 4096 tokens, so each step covers about 4 million tokens. We train for a total of 1.8 trillion tokens.
3.3. Training Dataset and Dataloader
@@
Our pre-training dataset includes a mix of data from various publicly available sources. It does not contain any data from Amazon or AWS customers, products, or services. We designed a novel dataloader for online example packing to process our datasets for training. It performs both tokenization and packing online during training. In contrast to the offline packing approach, this saves a lot of developer time and compute resources specially for large datasets, and multiple experiments with different datasets and tokenizers. Note that the online tokenization has no impact on training throughput as the tokenization for future samples/batch happens during forward-backward pass of current samples/batch - we use CPU for tokenization and Trainiumdevice for training, so the computations are in parallel. The online sample packing dataloader requires one or more dataset files in Apache Arrow format (Arrow, 2020). At first, all samples in the dataset are shuffled randomly and split into several subsets according to the total data parallel ranks. Each data split is treated as an independent data stream. For training efficiency, we use sample concatenation i.e. if a sample is shorter than the maximum sequence length of the model, we concatenate it with the following sample(s) to curate a sequence with total length equal or more than maximum sequence length. Any left over tokens from current sample is used in the following sample of the batch. The samples within a sequence is concatenated with a special end of sentence token. This gives the model necessary information to infer that the text separated by end of sentence token are unrelated (Brown et al., 2020). Note that the concatenated samples may be from very different sources or can be of different formats (e.g. natural language and codes). Finally, each batch of samples are tokenized on the fly. This online dataloader also has the functionality to save the internal states to checkpoint which enables us to resume the states from loaded checkpoint and continue training, in case of unexpected failures.
3.4. Orchestration
@@
For model training, we utilize a cluster with 64 trn1.32xlarge instances (nodes) with totalling to 1024 AWS Trainium  accelerators. Accelerators within same instance are connected with NeuronLink. The nodes within the cluster are interconnected through Elastic Fabric Adapter (EFA) [^6] https://aws.amazon.com/hpc/efa/. EFA is a network interface with uniquely designed operating system that bypasses traditional hardware interfaces, significantly enhancing performance for inter-node communications, a critical factor for collectives operations in distributed training.
We manage the cluster orchestration using Amazon EKS [^7] https://aws.amazon.com/eks/, a fully-managed Kubernetes service. Amazon EKS simplifies the deployment of Kubernetes both on AWS and on-premises environments. It autonomously manages the availability and scalability of Kubernetes control plane nodes, which are essential for tasks such as container scheduling, application availability management, cluster data storage, and other fundamental operations.
3.5. Training Efficiency
@@
LLaMA (Touvron et al., 2023a) model uses the efficient implementation features for pre-training on GPUs, that include xformer library, activation checkpointing, model parallelism, and computation/communication overlapping, etc. Similar features are also supported by Trainium  and Neuron SDK, as well as some unique enhancement such as BF16 with stochastic rounding. Below, we list the key features and configurations used in our model pretraining to improve the efficiency.
@@
Model Parallelism: We adopt tensor parallelism (TP) to shard the model into 8 TP degrees, and sequence parallelism (SP) to shard the 4096 sequence length into 8 SP degrees. This sharding configuration is observed with the highest throughput.
@@
Selective Activation Checkpointing: We use selective activation checkpointing (Korthikanti et al., 2022) to improve the training efficiency. It has slightly higher memory cost as full activation checkpointing, but increases the overall training throughput.
BF16 with Stochastic Rounding: Pre-training with full precision (FP32) is inefficient for large LLMs, but generic half-precision training (BF16 or FP16) often has numerical stability issues (Micikevicius et al., 2017). AWS Trainium features BF16 with stochastic rounding (SR) (Gupta et al., 2015) which is used in our training. Stochastic rounding, which theoretically provides an unbiased estimate of the input, prevents the computation precision-loss in BF16 by performing the rounding operations in a probabilistic manner. The chances of rounding-up/down are determined according to the relative distance from the two nearest representable values. This allows for small increments to accumulate over time, even when added to numbers of significantly higher magnitude, which leads to preciser results in gradient update. Empirically, we found that BF16 with SR shows the same convergence behavior as mixed precision training (Micikevicius et al., 2017) for HLAT, with higher training throughput and lower memory footprint.
Coalescing Layers with Same Inputs: We coalesced linear layers with the same inputs to reduce the communication in tensor and sequence parallelism, and increase efficiency of matrix operations. Specifically, the Q,K,V layers in an attention block are coalesced, and the two linear projections layers in SwiGLU (Shazeer, 2020) are also coalesced.
@@
Constant Attention Mask: As a decoder-only model, HLAT  pre-training uses a constant attention mask (lower-triangular) matrix. Instead of passing attention mask as an input tensor in model training, AWS Trainium  supports creating attention masks on Neuron Cores directly before use. This saves host memory usage, avoids redundant computation, and increases training throughput. To enable this feature in training script, the attention mask tensor is directly defined in attention block using torch.triu function and mapped to device=’xla’. The Neuron compiler will therefore enable constant attention mask optimization during compilation.
@@
--distribution-strategy=llm-training to enable the compiler to perform optimizations applicable to LLM training runs that shard parameters, gradients, and optimizer states across data-parallel workers. We also use --model-type=transformer that performs optimizations specific to transformer models. We set Neuron environment variable NEURON_FUSE_SOFTMAX=1 to enable compiler optimizations on custom lowering forSoftmax operation. Finally, we used NEURON_RT_ASYNC_EXEC_MAX_INFLIGHT_REQUESTS=3 to reduce training latency with asynchronous execution. This overlaps some executions of accelerators and host (CPU).
4. Training Process
4.1. Training Curves

### v2

@@ Abstract
Getting large language models (LLMs) to perform well on the downstream tasks requires pre-training over trillions of tokens. This typically demands a large number of powerful computational devices in addition to a stable distributed training framework to accelerate the training. The growing number of applications leveraging AI/ML led to a scarcity of the expensive conventional accelerators (such as GPUs), which emphasizes the need for the alternative specialized-accelerators that are scalable and cost-efficient. AWS Trainium is the second-generation machine learning accelerator purposely built for training large deep learning models. However, training LLMs with billions of parameters on AWS Trainium is challenging due to its relatively nascent software ecosystem. In this paper, we showcase HLAT: a family of 7B and 70B decoder-only LLMs pre-trained using 4096 AWS Trainium accelerators over 1.8 trillion tokens. The performance of HLAT is benchmarked against popular open source models including LLaMA and OpenLLaMA, which have been trained on NVIDIA GPUs and Google TPUs, respectively. On various evaluation tasks, we show that HLAT achieves model quality on par with the baselines of similar model size. We also open-source all the training scripts and configurations of HLAT [^1] https://github.com/awslabs/HLAT and share the best practice of using the NeuronX Distributed Training (NxDT), a customized distributed training library for AWS Trainium. Our work demonstrates that AWS Trainium powered by NxDT is able to successfully pre-train state-of-the-art LLM models with high performance and cost-effectiveness.
I Introduction
@@
Large language models (LLMs), based on transformer architecture [1] and trained on massive text data, are the most recent breakthrough in artificial intelligence. They not only show remarkable capabilities in understanding and generating text [2], but offer immense potential across diverse downstream tasks, such as machine translation [3], information retrieval [4], code generation [5] and so on [6].
@@
Pre-training is the crucial first step in building LLMs because it lays the foundation for their impressive capabilities. It initializes the model with random weights, and trains the model to convergence using tokens from a large text corpus. The training process is designed to be self-supervised. For decoder-only models, such as GPT [7] and LLaMA [8, 9, 10], the model is trained to predict the next token in a given sequence. Eventually, the model learns everything ranging from syntax and semantics to world knowledge and commonsense reasoning with a large amount of training data. Pre-training provides the raw material - the language skills and understanding, which facilitates subsequent fine-tuning for various downstream tasks.
@@
Since pre-training requires a large amount of training data (trillions of tokens), it demands highly on computational resources. Advanced AI accelerators, such as AWS Trainium [^2] https://aws.amazon.com/machine-learning/trainium, Google TPU, and NVIDIA A100/H100 GPUs, have been specifically designed for such workloads. These AI accelerators are often integrated with dedicated tensor processing units which offer fast matrix operations and high training throughput. They also have much larger on-chip memory (tens of GBs per accelerator) and high communication bandwidth (hundreds of Gbps) between accelerators across different machines, which allows pre-training of larger models with efficient hardware utilization.
@@
Even with the powerful AI accelerators, due to the sheer size and complexity of LLMs, it’s impractical to train them on a single device which has limited memory and processing power to handle the massive datasets, model parameters, and intricate calculations involved in LLM training. Practitioners rely on distributed training libraries [11, 12, 13, 14] to orchestrate a number of accelerators to conduct the training together. Distributed libraries can shard the model parameters and optimizer states across multiple accelerators with different kinds of parallelism strategy, allowing training of models with multi-billion parameters. They also spread the workload across multiple machines, effectively tapping into a combined pool of resources, to significantly reduce the training time.
@@
Although there have been many successful demonstrations of pre-training LLMs on conventional accelerators (GPUs and TPUs) using state-of-the-art distributed training libraries [11, 12, 13, 14], training LLMs with billions of parameters on AWS Trainium is still challenging. First, Trainium uses a relatively nascent software ecosystem ranging from runtime, compiler, to distributed training library. The training script developed for other accelerators needs to be adjusted to comply with the low-level APIs and operators supported by Trainium. Second, the optimal training configurations that ensure stable convergence and optimal training throughput may also differ from other accelerators, such as level of precision, dimensions of 3D parallelism, compiler flags and so on. On the other hand, Amazon EC2 trn1 instance, equipped with AWS Trainium accelerators, provides the comparable computation power to Amazon EC2 p4d instance, equipped with Nvidia A100 40GB GPUs, but comes with only $\sim$60of the price. This makes it appealing to fully utilize the compute power of AWS Trainium for LLM pre-training.
In this paper, we make the following contributions:
@@
  • Our work is the first to reproduce a SOTA LLM model on a completely new hardware - AWS Trainium. We end-to-end pre-train HLAT-7B and HLAT-70B (High-quality LLM pre-trained on AWS Trainium), following the architecture described in [9] from scratch. The pre-training covers 1.8 trillion tokens and is performed on up to 256 Amazon EC2 trn1.32xlarge instances with totalling up to 4096 AWS Trainium accelerators.
@@
  • We evaluate and show that both HLAT-7B and HLAT-70B perform comparable to models of similar size trained on other AI accelerators including LLaMA and OpenLLaMA. The evaluation is performed over a variety of tasks including commonsense reasoning, world knowledge, MMLU[15], math, coding, etc.
  • We propose multiple techniques to improve training efficiency on AWS Trainium such as a novel online dataloader, layer coalesing, selective activation checkpointing, precision strategy, and fault recovery mechanisms. These techniques save significant time and compute resources for pretraining on large datasets, and can be applied to other accelerators (GPU, TPU, etc.) as well.
  • We open-source all the training scripts and configurations of HLAT-7B and HLAT-70B models [^3] https://github.com/awslabs/HLAT, including the Pytorch training script, model definition script, and training configuration files (including all hyper-parameters). We also share best practices of pre-training on AWS Trainium and NeuronX Distributed Training (NxDT), such as sharding strategies, training precisions, compiler settings, etc. With those artifacts, practitioners can easily reproduce HLAT and pre-train their own custom LLMs on AWS Trainium.
II Background - Distributed Training on AWS Trainium
@@
AWS Trainium is the second-generation machine learning accelerator that AWS purposely built for deep learning training. Each Trainium accelerator includes two NeuronCores. Each NeuronCore has 16 GB of high-bandwidth memory, and delivers up to 95 TFLOPS of FP16/BF16 compute power. In this study, we trained our model on Amazon EC2 trn1.32xlarge instances: each instance is equipped with 16 Trainium accelerators, and supports 800 Gbps intra-instance network bandwidth through NeuronLink. The aggregating compute power of Amazon EC2 trn1.32xlarge is 3040 TFLOPS in FP16/BF16, slightly higher to its GPU instance counterpart Amazon EC2 p4d.24xlarge at 2496 TFLOPS, but at a much lower price (trn1.32xlarge \$21.50 vs. p4d.24xlarge \$32.77).
@@
AWS Neuron is a software development kit (SDK) [^4] https://github.com/aws-neuron/aws-neuron-sdk with a compiler, runtime, and profiling tools that unlocks high-performance and cost-effective deep learning acceleration on AWS Trainium. Neuron is natively integrated with PyTorch [14] and TensorFlow [16], and offers features such as FP32 autocasting, stochastic rounding, collective communication, custom operators, and so on.
NeuronX Distributed Training (NxDT) [^5] https://awsdocs-neuron.readthedocs-hosted.com/en/latest/libraries/nxd-training/overview.html, as part of Neuron SDK, is developed to enable high-efficiency distributed training on Trainium: NxDT supports a variety of distributed training techniques, such as 3D parallelism [12], i.e., Tensor Parallelism (TP), Pipeline Parallelism (PP) and Data Parallelism (DP). To reduce the activation memory during training, activation checkpointing [17] and sequence parallelism [18] are naturally supported with the 3D parallelism. NxDT also supports Zero Redundancy Optimizer Stage 1 (ZeRO-1) [19] to shard optimizer states, which can be applied simultaneously with 3D parallelism. NxDT provides unified interfaces to port custom models and run training scripts on AWS Trainium. To train models from Huggingface transformers library on Trainium accelerators with NxDT, it only requires simple code changes in certain layers of the model. NxDT supports TP with mixed degrees, i.e., users can use more than one TP degrees to shard different model parameters. This is helpful when some model parts of LLMs are not compatible with a unified large TP degree, e.g., with Grouped Query Attention (GQA) [20]. Finally, NxDT supports automatic fault recovery and checkpointing. In case of hardware failures or communication timeouts, NxDT can automatically restart training from latest auto-saved checkpoints without manual intervention, which is critical for maintaining system uptime and training efficiency.
III Method
III-A Model Architecture and Hyperparameters
@@
HLAT models adopt the decoder-only transformer architecture and apply same modifications used in LLaMA [8, 9, 10], including pre-normalization with RMSNorm, SwiGLU activation function, and Rotary Embeddings. HLAT-70B in addition applies GQA [20] with group size of 8. The models are trained with a maximum sequence length of 4096.
@@
We adopt training hyperparameters used in LLaMA2 [9]. Specifically, the global batch size is 1024 sequences, so each step covers about 4 million tokens. We use a cosine learning rate scheduler. The maximum learning rate is $3e^{-4}$ for HLAT-7B, and $1.5e^{-4}$ for HLAT-70B. The minimum learning rate decays to $10\%$ of maximum learning rate. We use a linear warmup of 2000 steps. The overall learning rate scheduler is plotted in Figure 1. We use AdamW optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.95$. We use weight decay value of 0.1 for all parameters, including normalization weights. Gradient-norm clipping of 1.0 is applied for training stability.
III-B Training Dataset and Dataloader
Our pre-training dataset includes RedPajama-1T [21], peS2o [22], and OpenWebMath [23]. HLAT-7B is purely trained on RedPajama-1T with 1.8 trillion tokens. HLAT-70B is initially trained on RedPajama-1T over 1.4 trillion tokens and is then continually trained on an up-sampled dataset with RedPajama-1T, peS2o and OpenWebMath for 400B tokens (see Section V-C for details).
@@ III-B Training Dataset and Dataloader
We designed a novel dataloader which performs both tokenization and packing online during training. The dataloader takes one or more dataset files in Apache Arrow format [24]. All samples are randomly shuffled and split into several subsets according to the total DP ranks. Each data split is treated as an independent data stream. For training efficiency, we use sample concatenation, i.e., if a sample is shorter than the maximum sequence length of the model, we concatenate it with the following sample(s) to curate a sequence with total length equal or more than maximum sequence length. Any left over tokens from current concatenated sequence is used in the following sequence. The samples within a concatenated sequence is concatenated with a special end of sentence (EOS) token. This gives the model necessary information to infer that the text separated by EOS token are unrelated [7]. Note that samples in the same concatenated sequence may be from very different sources or can be of different formats (e.g. natural language and codes). Finally, each batch of sequences are tokenized on the fly during the training.
The online tokenization has no impact on training throughput as the tokenization for future samples/batch happens during forward-backward pass of current samples/batch - we use CPU for tokenization and Trainium devices for training, so the computations are in parallel. In comparison, the offline dataloader requires pre-tokenization of entire datasets which costs a lot of developer time and compute resources for large datasets. The offline Nemo dataloader[25], for example, takes about 154 hours to pre-tokenize a dataset with 2 trillion tokens.
III-C Orchestration
@@
HLAT pre-training is performed on clusters with 256 trn1.32xlarge instances (nodes), totalling to 4096 AWS Trainium accelerators. Both Amazon EKS and AWS ParallelCluster can effectively manage the training cluster. Accelerators within same node are connected with NeuronLink [^6] https://awsdocs-neuron.readthedocs-hosted.com/en/latest/general/arch/neuron-hardware/trainium.html. The nodes within the cluster are interconnected through Elastic Fabric Adapter (EFA) [^7] https://aws.amazon.com/hpc/efa/. EFA is a network interface with uniquely designed operating system that bypasses traditional hardware interfaces, significantly enhancing performance for inter-node communications, a critical factor for collectives operations in distributed training.
III-D Training Efficiency
@@
LLaMA [8] model uses the efficient implementation features for pre-training on GPUs, that include xformer library, activation checkpointing, model parallelism, and computation/communication overlapping, etc. Similar features are also supported by Trainium and Neuron SDK, as well as some unique enhancement such as BF16 with stochastic rounding. Below, we list the key features and configurations used in our model pretraining to improve the efficiency.
@@
Model Parallelism: HLAT-7B is pre-trained over 64 nodes with TP=8, PP=1, and DP=256. HLAT-70B is pre-trained over 256 nodes with TP=32, PP=8, and DP=32. Both use sequence parallel (SP). This sharding configuration is observed to have the highest throughput.
@@
Selective Activation Checkpointing: We use selective activation checkpointing [18] to improve the training efficiency. It has slightly higher memory cost as full activation checkpointing, but increases the overall training throughput.
Training precision: Pre-training with full precision (FP32) is in-efficient for large LLMs, but generic half-precision training (BF16 or FP16) often has numerical stability issues [26]. On GPU, mixed precision training [26] is widely used to achieve similar precision as full FP32 with better efficiency. For HLAT-7B, we used BF16 with stochastic rounding (SR) [27], featured by AWS Trainium. Stochastic rounding, which theoretically provides an unbiased estimate of the input, prevents the computation precision-loss in BF16 by performing the rounding operations in a probabilistic manner. Empirically, we found that BF16 with SR shows the same convergence behavior as mixed precision training for HLAT-7B, with higher training throughput and lower memory footprint. However, on HLAT-70B, BF16 with SR shows worse convergence than mixed precision training. We found that mainly due to the nondeterministic normalization error introduced by stochastic rounding (see Section IV-C), and low-precision computations errors in gradient and attention operations. Such errors accumulate and signify on large models with more parameters and deeper layers. Therefore, we developed a standard mixed precision training strategy. Specifically, unless specified, all computation and storage use BF16 without stochastic rounding. This strategy uses FP32 in precision sensitive operators, local gradient accumulation, and global gradient synchronization. It also uses master weights and FP32 optimizer states. Finally, it uses ZeRO-1 [11] for memory efficiency.
@@
Constant Attention Mask: As a decoder-only model, HLAT pre-training uses a constant attention mask (lower-triangular) matrix. Instead of passing attention mask as an input tensor in model training, AWS Trainium supports creating attention masks on accelerators directly before use. The masked-out tensors will not be computed in the first place, which avoids redundant computation, saves host memory usage, and increases training throughput. To enable this feature in training script, the attention mask tensor is directly defined using torch.triu function and mapped to device=’xla’. The Neuron compiler will therefore enable the optimization during compilation.
Coalescing Layers with Same Inputs: We coalesced linear layers with the same inputs to reduce the communication in tensor and sequence parallelism, and increase efficiency of matrix operations. Specifically, the Q,K,V layers in an attention block are coalesced, and the two linear projections layers in SwiGLU [28] are also coalesced. Listing 1 shows a code snippet for implementation. Note that this technique also applies to other accelerators such as GPU and TPU.
1# Without layer coalescing
2query_states = q_proj(hidden_states) # hidden_size
3key_states = k_proj(hidden_states) # hidden_size
4value_states = v_proj(hidden_states) # hidden_size
5
6# With layer coalescing
7qkv_states = qkv_proj(hidden_states) # 3*hidden_size
8query_states, key_states, value_states = qkv_states.split(3, dim=2)
Listing 1: Example of layer coalescing.
@@
--distribution-strategy=llm-training and --model-type=transformer to enable the compiler to perform optimizations applicable to LLM (transformer model) training runs that shard parameters, gradients, and optimizer states across data-parallel workers. We set Neuron environment variable NEURON_FUSE_SOFTMAX=1 to enable compiler optimizations on custom lowering for Softmax operation. Finally, we used NEURON_RT_ASYNC_EXEC_MAX_INFLIGHT_REQUESTS=3 to reduce training latency with asynchronous execution. This overlaps some executions of accelerators and host (CPU).
IV Training Process
IV-A Training Curves